# Decoding Translation-Related Functional Sequences in 5 UTRs Using Interpretable Deep Learning Models

### v1

@@ Introduction
Table 1 demonstrates that UTR-STCNet consistently achieves state-of-the-art performance across the three 5’UTR translation datasets: MPRA-U, MPRA-H, and MPRA-V. The model is comprehensively evaluated in terms of its generalization capability and predictive accuracy. On the MPRA-U dataset, UTR-STCNet achieves approximately 11.4% lower prediction error than UTR-LM, reflecting superior modeling accuracy.This result suggests that the model more effectively captures the complex nonlinear relationships between sequence features and expression levels, maintaining superior predictive performance even for sequences with high expression or strong regulatory signals. On the MPRA-H dataset, UTR-STCNet outperforms mRNA2vec by around 4.8%, highlighting its enhanced ability to model gene expression on real human-derived sequences. This performance gain demonstrates that UTR-STCNet is better equipped to learn intrinsic regulatory features and expression patterns in biologically authentic contexts, providing improved reliability and application potential in human transcriptome modeling. On the MPRA-V dataset, UTR-STCNet achieves a 3.1% improvement in Spearman correlation compared to MTrans. This dataset includes variable-length sequences, designed to rigorously evaluate the model’s robustness to input heterogeneity and its ability to capture long-range dependencies. The observed performance further validates UTR-STCNet’s architectural advantages in modeling complex contextual structures and its practical applicability to diverse 5’UTR length distributions found in real biological scenarios.
@@
To evaluate the contribution of each core module in UTR-STCNet, we perform an ablation study on the MPRA-H dataset, as shown in Table 2. Specifically, we examine three ablation variants: removing both the Saliency-Aware Token Clustering Module STCM and the Saliency-Guided Aggregation Module SGA, removing only SGA while retaining STCM, and removing only STCM while keeping SGA. The results indicate that the combination of STCM and SGA yields the strongest performance, with an $R^{2}$ score of 0.849, Spearman correlation of 0.914, and RMSE of 0.471. In comparison, removing the SGA module alone reduces the $R^{2}$ to 0.837, showing that SGA plays a crucial role in preserving local continuity and guiding attention-weighted channel fusion. Similarly, removing STCM reduces the $R^{2}$ to 0.833, confirming that saliency-aware token clustering contributes to compressing redundant regions and enhancing structural coherence. When both modules are removed, the model performs worst, with $R^{2}$ dropping to 0.826 and RMSE increasing to 0.533. This highlights the limitations of directly applying vanilla Transformers to long UTR sequences without guided compression or context-aware aggregation. These findings validate the effectiveness of both modules. STCM improves input compactness through structural guidance, while SGA enhances downstream predictions by emphasizing important signals during aggregation. Together, they contribute synergistically to the model’s overall accuracy and generalization.
Biological Relevance of Model-Identified Regulatory Features
@@
Impact of Adjacent Nucleotides to TG on MRL Prediction.
@@
We analyzed the distribution of experimen tally measured MRL values for different TG-associated nu cleotide combinations. As shown in Figure 4a, sequences containing the ATG trinucleotide exhibited significantly lower MRL values compared to other combinations, suggesting that ATG may act as a translation-suppressive element (Andreev et al. 2022). In contrast, varying the nucleotide immediately following TG (A, C, G, or T) showed no substantial differences in MRL distributions Figure 4b, indicating that among trinucleotides containing TG, the presence of upstream ATG (uATG) is specifically associated with a negative regulatory effect (Tsotakos et al. 2015).
To validate the negative regulatory effect of ATG on MRL, we conducted an extreme test by replacing any non-A base immediately preceding each TG with A, while leaving other bases unchanged. This resulted in a significant overall decrease in MRL. This resulted in a significant overall decrease in MRL Figure 5a, further supporting the conclusion that ATG negatively impacts translation efficiency. Finally, to confirm whether the inhibitory effect is caused by the ATG triplet or the TG sequence itself, we randomly replaced all TGs with non-TG combinations Figure 5b. The prediction results indicated that while TG slightly reduced reaction efficiency, the primary inhibitory effect was attributed to the ATG sequence.
@@
To evaluate whether the model can capture regulatory patterns near translation initiation sites, we analyzed the nucleotide composition surrounding high-saliency TG motifs. For each TG site within high-scoring regions, we extracted four upstream and one downstream nucleotide and computed the positional nucleotide frequency distribution. The analysis revealed that the surrounding sequence context closely resembles the classical Kozak motif Figure 6 (Kozak 1986), indicating that the model effectively captures known regulatory features associated with translation efficiency.
Conclusion
In this work, we propose UTR-STCNet, a deep learning framework for modeling translational regulation from 5’UTR sequences. By combining a Saliency-Aware Token Clustering (SATC) module with a Saliency-Guided Transformer (SGT) block, UTR-STCNet captures long-range dependencies and key regulatory motifs with high interpretability. Extensive experiments demonstrate that the model consistently outperforms existing approaches across species and cell types, achieving state-of-the-art performance in mean ribosome load (MRL) prediction. In addition to accurate inference, UTR-STCNet reveals biologically meaningful patterns, recovering canonical elements such as uAUGs and Kozak motifs. These results highlight UTR-STCNet as a promising tool for advancing mRNA design and understanding translation control.

### v2

@@ IV-C Performance Comparison
Table I demonstrates that UTR-STCNet consistently achieves state-of-the-art performance across the three 5’UTRs translation datasets: MPRA-U, MPRA-H, and MPRA-V. The model is comprehensively evaluated in terms of its generalization capability and predictive accuracy. On the MPRA-U dataset, UTR-STCNet achieves approximately 11.4% lower prediction error than UTR-LM, reflecting superior modeling accuracy.This result suggests that the model more effectively captures the complex nonlinear relationships between sequence features and expression levels, maintaining superior predictive performance even for sequences with high expression or strong regulatory signals. On the MPRA-H dataset, UTR-STCNet outperforms mRNA2vec by around 4.8%, highlighting its enhanced ability to model gene expression on real human-derived sequences. This performance gain demonstrates that UTR-STCNet is better equipped to learn intrinsic regulatory features and expression patterns in biologically authentic contexts, providing improved reliability and application potential in human transcriptome modeling. On the MPRA-V dataset, UTR-STCNet achieves a 3.1% improvement in Spearman correlation compared to MTrans. This dataset includes variable-length sequences, designed to rigorously evaluate the model’s robustness to input heterogeneity and its ability to capture long-range dependencies. The observed performance further validates UTR-STCNet’s architectural advantages in modeling complex contextual structures and its practical applicability to diverse 5’UTRs length distributions found in real biological scenarios.
@@
To evaluate the contribution of each core module in UTR-STCNet, we perform an ablation study on the MPRA-H dataset, as shown in Table II. Specifically, we examine three ablation variants: removing both the STCM and the SGA, removing only SGA while retaining STCM, and removing only STCM while keeping SGA. The results indicate that the combination of STCM and SGA yields the strongest performance, with an $R^{2}$ score of 0.849, Spearman correlation of 0.914, and RMSE of 0.471. In comparison, removing the SGA module alone reduces the $R^{2}$ to 0.837, showing that SGA plays a crucial role in preserving local continuity and guiding attention-weighted channel fusion. Similarly, removing STCM reduces the $R^{2}$ to 0.833, confirming that saliency-aware token clustering contributes to compressing redundant regions and enhancing structural coherence. When both modules are removed, the model performs worst, with $R^{2}$ dropping to 0.826 and RMSE increasing to 0.533. This highlights the limitations of directly applying vanilla Transformers to long UTRs sequences without guided compression or context-aware aggregation. These findings validate the effectiveness of both modules. STCM improves input compactness through structural guidance, while SGA enhances downstream predictions by emphasizing important signals during aggregation. Together, they contribute synergistically to the model’s overall accuracy and generalization.
IV-D Biological Relevance of Model-Identified Regulatory Features
@@
IV-E Impact of Adjacent Nucleotides to TG on MRL Prediction.
@@
We analyzed the distribution of experimentally measured MRL values for different TG-associated nucleotide combinations. As shown in Figure 4a, sequences containing the ATG trinucleotide exhibited significantly lower MRL values compared to other combinations, suggesting that ATG may act as a translation-suppressive element [3]. In contrast, varying the nucleotide immediately following TG (A, C, G, or T) showed no substantial differences in MRL distributions Figure 4b, indicating that among trinucleotides containing TG, the presence of upstream ATG (uATG) is specifically associated with a negative regulatory effect [23].
To validate the negative regulatory effect of ATG on MRL, we conducted an extreme test by replacing any non-A base immediately preceding each TG with A, while leaving other bases unchanged. This resulted in a significant overall decrease in MRL Figure 5a, further supporting the conclusion that ATG negatively impacts translation efficiency. Finally, to confirm whether the inhibitory effect is caused by the ATG triplet or the TG sequence itself, we randomly replaced all TGs with non-TG combinations Figure 5b. The prediction results indicated that while TG slightly reduced reaction efficiency, the primary inhibitory effect was attributed to the ATG sequence.
@@
To evaluate whether the model can capture regulatory patterns near translation initiation sites, we analyzed the nucleotide composition surrounding high-saliency TG motifs. For each TG site within high-scoring regions, we extracted four upstream and one downstream nucleotide and computed the positional nucleotide frequency distribution. The analysis revealed that the surrounding sequence context closely resembles the classical Kozak motif Figure 6 [17], indicating that the model effectively captures known regulatory features associated with translation efficiency.
V Conclusion
In this work, we propose UTR-STCNet, a deep learning framework for modeling translational regulation from 5’UTRs sequences. By combining a Saliency-Aware Token Clustering module with a Saliency-Guided Transformer block, UTR-STCNet captures long-range dependencies and key regulatory motifs with high interpretability. Extensive experiments demonstrate that the model consistently outperforms existing approaches across species and cell types, achieving state-of-the-art performance in MRL prediction. In addition to accurate inference, UTR-STCNet reveals biologically meaningful patterns, recovering canonical elements such as uAUGs and Kozak motifs. These results highlight UTR-STCNet as a promising tool for advancing mRNA design and understanding translation control.
Acknowledgment
We are grateful to the members of the Yu Laboratory and many colleagues for their valuable comments and suggestions. This work was supported by the Major Project of Guangzhou National Laboratory (Grant No.GZNL2024A01003), the National Natural Science Foundation of China (Grant No.32470634), and the Guangdong Basic and Applied Basic Research Foundation (Grant No.2024B1515020080).